OpenStudy (anonymous):

How do I change this number to standard scientific notation 5.408?

OpenStudy (anonymous):

@anitasonia98

OpenStudy (zzr0ck3r):

well, it is already almost there. You just need to write \(5.408\times 10^0\)

OpenStudy (anonymous):

really?

OpenStudy (zzr0ck3r):

yes, scientific notation is in the form \(a.bcde..\times 10^k\)

OpenStudy (zzr0ck3r):

in other words we only want one number before the decimal

OpenStudy (anonymous):

How do I write 72.41 x 10^-2 in standard scientific notaiton?

OpenStudy (zzr0ck3r):

if it were \(10.58\) then we would write \(1.058\times 10^{-1}\)

OpenStudy (anonymous):

Ooh Thanks

OpenStudy (zzr0ck3r):

we want one number in front of the decimal If you need to move the decimal to the left to make that happen, then you subtract one from the power of 10 if you need to move the decimal to the right then you add one to the power of 10

OpenStudy (freckles):

is it weird they must the word standard in front of scientific notation? seems like an extra word to me.

OpenStudy (zzr0ck3r):

I think standard means base 10?

OpenStudy (freckles):

so there is unstandard scientific notation? :p

OpenStudy (zzr0ck3r):

\(72.41\times 10^{-2}=7.241\times 10^{-3}\)

OpenStudy (zzr0ck3r):

I have no idea, but if there was a standard, surely it would be the one we know :)

OpenStudy (zzr0ck3r):

Maybe some Pigmy tribe uses something different.

OpenStudy (zzr0ck3r):

ps I have that backwards, if we move left add 1, if we move right subtract 1.